# Equivariant and Invariant Parametrized Topological Complexity

### v3

@@ 5. Computations for equivariant Fadell-Neuwirth fibrations
see [9, Equation 4.1] for the homotopy dimension of $F$. Since $d\geq 3$, the space $F\times F$ is simply connected. Consequently, the induced map $\pi_{1}(\overline{E}\times_{\overline{B}}\overline{E})\to\pi_{1}(\overline{B})$ is an isomorphism. This can be deduced from the long exact sequence of homotopy groups corresponding to fibration (8). This implies that every local coefficient on $\overline{E}\times_{\overline{B}}\overline{E}$ is a pullback of a local coefficient system on $\overline{B}$. As a result, we have $H^{k+1}(\overline{E}\times_{\overline{B}}\overline{E};\pi_{k}(*_{2t+s}\Omega F))=0$ for $k>(2t+s-1)(d-1)-1=(2t+s)(d-1)-d$. Thus, the obstruction classes vanishes for all $k$. ∎
We would like to thank Professor Mark Grant for suggesting the use of obstruction theory in the above proof to improve our earlier bound, and for explicitly explaining how to apply the cohomological dimension of the unordered configuration space in the Serre spectral sequence to show that the obstructions vanish.
We now turn our attention to the case $d=2$. For $d=2$, we note that the spaces $E$ and $B$ are aspherical, see [23, Lemma 3.4]. Since the maps $E\to\overline{E}$ and $B\to\overline{B}$ are covering maps, it follows that $\overline{E}$ and $\overline{B}$ are also aspherical, as $E\to\overline{E}$ and $B\to\overline{B}$ are covering maps. Therefore, we can apply the techniques developed by Grant in [22] to compute $\mathrm{TC}[\overline{p}\colon\overline{E}\to\overline{B}]$, as the map $\overline{p}\colon\overline{E}\to\overline{B}$ is a fibration of aspherical spaces with path-connected fibre $F$. More precisely, this is equivalent to computing the topological complexity of the group epimorphism $\overline{p}_{*}\colon\pi_{1}(\overline{E})\to\pi_{1}(\overline{B})$. First, we recall some definitions that will be useful for our discussion.
Definition 5.7 ([22, Definition 3.1]).
We say that a pointed map $f\colon X\to Y$ realizes the group homomorphism $\rho\colon G\to H$ if $X$ is a $K(G,1)$ space, $Y$ is a $K(H,1)$ space, and $f$ induces the homomorphism $\rho$ on fundamental groups.
Definition 5.8 ([13]).
@@
We note that a group of type $F$ is of type $FP$, see [13, Page 171].
Definition 5.9 ([13, Theorem 3.6]).
@@
We now use arguments presented in [24, Theorem 3] to prove the following lemma. This lemma, in a certain sense, generalizes part of [13, Theorem 4.3] from Poincaré duality groups to duality groups.
@@
Observe that all lattice points on the page $E_{2}$ that do not lie on the vertical line $q=h$ are zero. Since the differential $d_{r}$ maps $E_{r}^{p,q}$ to $E_{r}^{p+r,q-r+1}$, it follows that $H^{p}(K;H^{q}(H;\mathbb{Z}G))=E_{2}^{p,q}=E_{\infty}^{p,q}=H^{p+q}(G;\mathbb{Z}G)$. In particular, $H^{p}(K;\mathbb{Z}K)=0$ if and only if $H^{p+h}(G,\mathbb{Z}G)=0$. ∎
@@
Suppose $B_{s+t}$ is the braid group on $(s+t)$-strands. Then, by [23, Lemma 3.4], the fundamental group of $F(\mathbb{R}^{d},s+t)/\Sigma_{s}$ is
@@
where $\pi\colon B_{s+t}\to\Sigma_{s+t}$ is the canonical map. Moreover, by [23, Lemma 3.6], we have $B_{s+t}^{\Sigma_{s}}$ is a duality group of cohomological dimension $s+t-1$.
@@
by Theorem 4.29 and [22, Proposition 3.5].
If $Z$ is the centre of $B_{s+t}$, then $Z$ is infinite cyclic and the centre of $B_{s+t}^{\Sigma_{s}}$ is $Z$ as well, see [23, Lemma 3.7]. Hence, by [22, Theorem 3.5], it follows that
@@
$1\to\overline{P}_{t,s}\to B_{s+t}^{\Sigma_{s}}\times_{B_{s}}B_{s+t}^{\Sigma_{s}}\to B_{s+t}^{\Sigma_{s}}\to 1$
@@
$1\to\overline{P}_{t,s}\to\frac{B_{s+t}^{\Sigma_{s}}\times_{B_{s}}B_{s+t}^{\Sigma_{s}}}{\Delta(Z)}\to\frac{B_{s+t}^{\Sigma_{s}}}{Z}\to 1$ (10)
We note that $\overline{P}_{t,s}$ and $P_{s+t}$ are duality groups with $\mathrm{cd}(\overline{P}_{t,s})=t$ and $\mathrm{cd}(P_{s+t})=s+t-1$, see [23, Lemma 3.6].
We will now show that $B_{s+t}^{\Sigma_{s}}/Z$ is a duality group with $\mathrm{cd}(B_{s+t}^{\Sigma_{s}}/Z)=s+t-2$. As $t\geq 2$, the inclusion $Z\hookrightarrow B_{s+t}^{\Sigma_{s}}$ splits. This can be seen geometrically by the projection $B_{s+t}^{\Sigma_{s}}\to P_{2}$ obtained by forgetting the first $s+t-2$ strands, where $P_{2}$ is the pure braid group on $2$ strands. Hence,
@@
Note that the space $F(\mathbb{R}^{d},s+t)$ contains a finite $\mathrm{CW}$-complex $C$ which is a $\Sigma_{s+t}$-equivariant strong deformation retract, see [6, Theorem 3.13]. Hence, $C$ is also a $\Sigma_{s}$-equivariant strong deformation retract of $F(\mathbb{R}^{d},s+t)$. Hence, the homotopy equivalence $K(B_{s+t}^{\Sigma_{s}},1)\simeq_{h}F(\mathbb{R}^{d},s+t)/\Sigma_{s}\simeq_{h}C/\Sigma_{s}$ implies $K(B_{s+t}^{\Sigma_{s}},1)$ is homotopy equivalent to a finite $\mathrm{CW}$-complex. Therefore, the homotopy equivalence
@@
implies $K(B_{s+t}^{\Sigma_{s}}/Z,1)$ is homotopy equivalent to a finite $\mathrm{CW}$-complex, i.e., $B_{s+t}^{\Sigma_{s}}/Z$ is a group of type $F$. Then Lemma 5.10 applied to the extension $1\to Z\hookrightarrow B_{s+t}^{\Sigma_{s}}\to B_{s+t}^{\Sigma_{s}}/Z\to 1$ implies that $B_{s+t}^{\Sigma_{s}}/Z$ is a duality group of cohomological dimension $s+t-2$. Hence, by [4, Theorem 3.5], it follows that the middle group in (10) has cohomological dimension $s+2t-2$. ∎
@@
Note that when $t=1$, the group $B_{s+1}^{\Sigma_{s}}/Z$ is not torsion free, as shown in [23, Proposition 4.2]. Consequently, every $\mathrm{CW}$-complex of type $K(B_{s+1}^{\Sigma_{s}}/Z,1)$ is infinite dimensional. Therefore, the argument used in the preceding theorem fails for the case $t=1$.
@@
We expect the above conjecture to be true since the motion planning problem for two submarines constraint to the unknown position of $s$ mines (where the order in which mines are placed does not matter) should be more complex than of a single submarine. Hence,
@@
5.13 is true if $s=1$, see Remark 5.2 and [17, Theorem 1].
We note that we can get $\mathrm{TC}[\overline{F(\mathbb{R}^{2},s+1)}\to\overline{F(\mathbb{R}^{2},s)}]\leq 4+s$, using [8, Proposition 7.2] and Proposition 5.5 (3). Moreover, $\mathrm{TC}[\overline{F(\mathbb{R}^{2},s+1)}\to\overline{F(\mathbb{R}^{2},s)}]\geq 1+s$, by Theorem 5.3.
6. Acknowledgement
We thank Professor Mark Grant for his valuable feedback on an earlier version of this article, which significantly improved the paper in various aspects. In particular, we are grateful to him for suggesting an approach to establish a cohomological lower bound via the cohomology of orbit spaces, and for helping us prove Theorem 5.6 without any additional assumptions. We also thank Professor Jesus Gonzalez for pointing us to an appropriate reference concerning the homotopy dimension of the unordered configuration space. The first author would like to acknowledge IISER Pune - IDeaS Scholarship and Siemens-IISER Ph.D. fellowship for economical support. The second author acknowledges the support of NBHM through grant 0204/10/(16)/2023/R&D-II/2789.

### v4

@@ 5. Computations for equivariant Fadell-Neuwirth fibrations
see [PTCcolfree, Equation 4.1] for the homotopy dimension of $F$. Since $d\geq 3$, the space $F\times F$ is simply connected. Consequently, the induced map $\pi_{1}(\overline{E}\times_{\overline{B}}\overline{E})\to\pi_{1}(\overline{B})$ is an isomorphism. This can be deduced from the long exact sequence of homotopy groups corresponding to fibration (11). This implies that every local coefficient on $\overline{E}\times_{\overline{B}}\overline{E}$ is a pullback of a local coefficient system on $\overline{B}$. As a result, we have $H^{k+1}(\overline{E}\times_{\overline{B}}\overline{E};\pi_{k}(*_{2t+s}\Omega F))=0$ for k>(2t+s1)(d1)1=(2t+s)(d1)dk>(2t+s-1)(d-1)-1=(2t+s)(d-1)-d. Thus, the obstruction classes vanishes for all $k$. ∎
We now turn our attention to the case $d=2$. For $d=2$, we note that the spaces $E$ and $B$ are aspherical, see [grant2018artin-braid, Lemma 3.4]. Since the maps $E\to\overline{E}$ and $B\to\overline{B}$ are covering maps, it follows that $\overline{E}$ and $\overline{B}$ are also aspherical, as $E\to\overline{E}$ and $B\to\overline{B}$ are covering maps. Therefore, we can apply the techniques developed by Grant in [Grant-ParaTC-group] to compute $\mathrm{TC}[\overline{p}\colon\overline{E}\to\overline{B}]$, as the map $\overline{p}\colon\overline{E}\to\overline{B}$ is a fibration of aspherical spaces with path-connected fibre $F$. More precisely, this is equivalent to computing the topological complexity of the group epimorphism $\overline{p}_{*}\colon\pi_{1}(\overline{E})\to\pi_{1}(\overline{B})$. We begin by recalling some definitions that will be useful in the discussion. Recall that a pointed map $f\colon X\to Y$ is said to realize a group homomorphism $\rho\colon G\to H$ if $X$ is a $K(G,1)$ space, $Y$ is a $K(H,1)$ space, and $f$ induces the homomorphism $\rho$ on fundamental groups.
Definition 5.7 ([davis2000poincare]).
@@
We note that a group of type $F$ is of type $FP$, see [davis2000poincare, Page 171].
Definition 5.8 ([davis2000poincare, Theorem 3.6]).
@@
We now use arguments presented in [wall-poincare-duality, Theorem 3] to prove the following lemma. This lemma, in a certain sense, generalizes part of [davis2000poincare, Theorem 4.3] from Poincaré duality groups to duality groups.
@@
Since the differential $d_{r}$ maps $E_{r}^{p,q}$ to $E_{r}^{p+r,q-r+1}$, it follows that $H^{p}(K;H^{q}(H;\mathbb{Z}G))=E_{2}^{p,q}=E_{\infty}^{p,q}=H^{p+q}(G;\mathbb{Z}G)$. In particular, $H^{p}(K;\mathbb{Z}K)=0$ if and only if $H^{p+h}(G,\mathbb{Z}G)=0$. ∎
@@
Suppose $B_{s+t}$ is the braid group on $(s+t)$-strands. Then, by [grant2018artin-braid, Lemma 3.4], the fundamental group of $F(\mathbb{R}^{d},s+t)/\Sigma_{s}$ is
@@
where $\pi\colon B_{s+t}\to\Sigma_{s+t}$ is the canonical map. Moreover, by [grant2018artin-braid, Lemma 3.6], we have $B_{s+t}^{\Sigma_{s}}$ is a duality group of cohomological dimension s+t1s+t-1.
@@
by Theorem˜4.26 and [Grant-ParaTC-group, Proposition 3.5].
If $Z$ is the centre of $B_{s+t}$, then $Z$ is infinite cyclic and the centre of $B_{s+t}^{\Sigma_{s}}$ is $Z$ as well, see [grant2018artin-braid, Lemma 3.7]. Hence, by [Grant-ParaTC-group, Theorem 3.5], it follows that
@@
$1\to\overline{P}_{t,s}\to B_{s+t}^{\Sigma_{s}}\times_{B_{s}}B_{s+t}^{\Sigma_{s}}\to B_{s+t}^{\Sigma_{s}}\to 1.$
@@
$1\to\overline{P}_{t,s}\to\frac{B_{s+t}^{\Sigma_{s}}\times_{B_{s}}B_{s+t}^{\Sigma_{s}}}{\Delta(Z)}\to\frac{B_{s+t}^{\Sigma_{s}}}{Z}\to 1.$ (13)
We note that $\overline{P}_{t,s}$ and $P_{s+t}$ are duality groups with $\mathrm{cd}(\overline{P}_{t,s})=t$ and $\mathrm{cd}(P_{s+t})=s+t-1$, see [grant2018artin-braid, Lemma 3.6].
We will now show that $B_{s+t}^{\Sigma_{s}}/Z$ is a duality group with $\mathrm{cd}(B_{s+t}^{\Sigma_{s}}/Z)=s+t-2$. As $t\geq 2$, the inclusion $Z\hookrightarrow B_{s+t}^{\Sigma_{s}}$ splits. This can be seen geometrically by the projection $B_{s+t}^{\Sigma_{s}}\to P_{2}$ obtained by forgetting the first s+t2s+t-2 strands, where $P_{2}$ is the pure braid group on $2$ strands. Hence,
@@
Note that the space $F(\mathbb{R}^{d},s+t)$ contains a finite $\mathrm{CW}$-complex $C$ which is a $\Sigma_{s+t}$-equivariant strong deformation retract, see [B-Z, Theorem 3.13]. Hence, $C$ is also a $\Sigma_{s}$-equivariant strong deformation retract of $F(\mathbb{R}^{d},s+t)$. Hence, the homotopy equivalence $K(B_{s+t}^{\Sigma_{s}},1)\simeq_{h}F(\mathbb{R}^{d},s+t)/\Sigma_{s}\simeq_{h}C/\Sigma_{s}$ implies $K(B_{s+t}^{\Sigma_{s}},1)$ is homotopy equivalent to a finite $\mathrm{CW}$-complex. Therefore, the homotopy equivalence
@@
implies $K(B_{s+t}^{\Sigma_{s}}/Z,1)$ is homotopy equivalent to a finite $\mathrm{CW}$-complex, i.e., $B_{s+t}^{\Sigma_{s}}/Z$ is a group of type $F$. Then Lemma˜5.9 applied to the extension $1\to Z\hookrightarrow B_{s+t}^{\Sigma_{s}}\to B_{s+t}^{\Sigma_{s}}/Z\to 1$ implies that $B_{s+t}^{\Sigma_{s}}/Z$ is a duality group of cohomological dimension s+t2s+t-2. Hence, by [bieri-duality-groups, Theorem 3.5], it follows that the middle group in (13) has cohomological dimension s+2t2s+2t-2. ∎
@@
Note that when $t=1$, the group $B_{s+1}^{\Sigma_{s}}/Z$ is not torsion free, as shown in [grant2018artin-braid, Proposition 4.2]. Consequently, every $\mathrm{CW}$-complex of type $K(B_{s+1}^{\Sigma_{s}}/Z,1)$ is infinite dimensional. Therefore, the argument used in the preceding theorem fails for the case $t=1$.
@@
We expect the above conjecture to be true, since the motion planning problem for two submarines constrained by the unknown positions of $s$ mines (where the order in which the mines are placed does not matter) should be more complex than that for a single submarine. Hence,
@@
˜5.12 is true if $s=1$, see Remark˜5.2 and [TC-of-configuration-spaces, Theorem 1].
We note that we can get $\mathrm{TC}[\overline{F(\mathbb{R}^{2},s+1)}\to\overline{F(\mathbb{R}^{2},s)}]\leq 4+s$, using [farber-para-tc, Proposition 7.2] and Proposition˜5.5. Moreover, $\mathrm{TC}[\overline{F(\mathbb{R}^{2},s+1)}\to\overline{F(\mathbb{R}^{2},s)}]\geq 1+s$, by Theorem˜5.3.
6. Acknowledgements
We thank Professor Mark Grant for his valuable feedback on an earlier version of this article, which significantly improved the paper in various aspects. In particular, we are grateful to him for suggesting an approach to establish a cohomological lower bound via the cohomology of orbit spaces, for recommending the use of obstruction theory in Theorem˜5.6 to improve our earlier bound, and for clearly explaining how to apply the cohomological dimension of the unordered configuration space in the Serre spectral sequence to show that the obstructions vanish. We also thank Professor Jesus Gonzalez for pointing us to an appropriate reference on the equivariant $\mathrm{CW}$-complex structure on the ordered configuration space, from which we can determine the homotopy dimension of the unordered configuration space. The first author would like to acknowledge IISER Pune - IDeaS Scholarship and Siemens-IISER Ph.D. fellowship for economical support. The second author acknowledges the support of NBHM through grant 0204/10/(16)/2023/R&D-II/2789 and DST–INSPIRE Faculty Fellowship (Faculty Registration No. IFA24-MA218), Department of Science and Technology, Government of India.